OpenStudy (anonymous):

The density of a fish tank is 0.2fish over feet cubed. There are 8 fish in the tank. What is the volume of the tank?

OpenStudy (anonymous):

This is a unit conversion problem; although it might not look like it. \[\frac{8 fish }{ m^3 }\]=\[\frac{m^3 }{8 fish}\] What do we do when we want to get rid of a fractions denominator? We multiply by the denominator, or 0.2 fish =\[\frac{m^3 }{8 fish}\]*0.2 fish The two "fish" units cancel out, leaving us with 8/0.2 cubic meters, or 40 cubic meters.

OpenStudy (anonymous):

I see the original problem mentioned feet, while I said meters by accident; the same method and answer applies. The key here is realizing that you can treat units just like numbers.

OpenStudy (anonymous):

Is there like a specific formula for these type of problems?

OpenStudy (anonymous):

No, there's not. You wouldn't want to use it even if there was one. The idea is you multiply by whatever the problem gives you until you get out the units you want. In this case, we multiplied by 0.2 fish because we wanted to get rid of fish on the bottom. An easy way to remember is by trying it with miles per hour. What does per mean in math terms? Divided by. So, if I'm going 40 miles per hour, I'm going 40 miles/hour How long have I driven in 4 hours? \[\frac{ 40 miles }{ hour }\]* 4 hours The hours cancel out, leaving us with 4* 40 miles, or 160 miles. This means we will have traveled 160 miles in 4 hours. You probably did this all in your head as soon as you read it; just realize that's what you're doing in this type of problems, except instead of miles per hour you have fish per cubic foot.

OpenStudy (anonymous):

Awesome! I didn't think of it that way! Thank you so much!!!!

OpenStudy (anonymous):

Actually, the answer is 40 ft^3 0.2 fish per feet cubed, so it's simple just 8 x 5 which adds up to 40ft. I just took the test a few minutes ago and 40 was the answer. :)